OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

D = distance, S = speed, T = time, A = car A, B = car b \[D_A=D_B\] \[S_A+15=S_B\] \[T_A =2\] \[T_B=1.5\] And remember... \[S = \frac{ D }{ T }\] Can you do the rest?

OpenStudy (anonymous):

@StudyGurl14 How do you get speed A? All I know is it was 15 mph slower than car B

OpenStudy (studygurl14):

\[S_A=\frac{ D_A }{ T_A} ; S_B = \frac{D_B}{T_B}\] Does that help?

OpenStudy (anonymous):

So the distance divided by time is the speed of car A, but i dont know the distance.

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

Because they travel the same distance, you can just use D instead of D_A and D_B... \[S_A=\frac{D}{T_A}\] \[S_B = \frac{D}{T_B}\] \[D=S_A \times T_A ; D = S_B \times T_B\] That should help.

OpenStudy (anonymous):

I'm still very confused about this question. I don't know speed or distance. all I know is time and that B was 15 mph faster. How do you divide a letter?

OpenStudy (anonymous):

@StudyGurl14 Is it 60?

OpenStudy (studygurl14):

YES! OMG great job!

OpenStudy (studygurl14):

do you know how to do it, or did you find that on the internet? do you want me to show you the right way to do it?

OpenStudy (anonymous):

@StudyGurl14 I failed the test, and it said that was the correct answer. I thought it was 50

OpenStudy (anonymous):

i got 2 questions wrong on the test :S

OpenStudy (anonymous):

@StudyGurl14 Can you show me the right way to do it? I need to know to redo the test.

OpenStudy (studygurl14):

two questions wrong isn't failing...

OpenStudy (anonymous):

when there are 5 questions it is.

OpenStudy (studygurl14):

oh...then it is a quick check. who cares. those are only worht 5% of grade

OpenStudy (studygurl14):

do you want me to show you how?

OpenStudy (anonymous):

it was an assesment, yes please

OpenStudy (studygurl14):

okay. hold on.

OpenStudy (studygurl14):

You have all this information: \[S_A=\frac{D_A}{T_A} ; S_B=\frac{D_B}{T_B}\] \[D_A=S_A \times T_A ; D_B=S_B \times T_B\] \[D_A=D_B\] so... \[D = S_A \times T_A ; D=S_B \times T_B\] so... \[S_A \times T_A = S_B \times T_B\] And we know... \[S_A+15=S_B\] \[T_A = 2 ; T_B = 1.5\] Rearrange: \[S_A = S_B - 15\] Substitute: \[(S_B-15) \times (2) = S_B \times (1.5)\] Solve for S_B (which is the speed of car B) Can you do the rest yourself?

OpenStudy (anonymous):

@StudyGurl14 it is like you are speaking a foreign language. I don't understand it.

OpenStudy (studygurl14):

What don't you understand? Be specific

OpenStudy (anonymous):

How you got (SB−15)×(2)=SB×(1.5)

OpenStudy (anonymous):

or how to solve

OpenStudy (anonymous):

(Sb-15)x(2)=Sbx(1.5)*

OpenStudy (studygurl14):

Okay, let me finish it for you.... \[(S_B-15)\times (2) = S_B \times (1.5)\] \[2S_B - 30 = 1.5S_B\] \[-30 = -0.5S_B\] \[S_B = 60\]

OpenStudy (studygurl14):

still confused?

OpenStudy (anonymous):

Ohhhh

OpenStudy (anonymous):

I have this one, if you want me to open up another question so you get another medal i will Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 4 h, and Car B traveled the distance in 4.5 h. Car B traveled 5 mph slower than Car A. How fast did Car B travel?

OpenStudy (anonymous):

How do you get the equation?

OpenStudy (studygurl14):

open new question please

OpenStudy (studygurl14):

how did i get what equation?

OpenStudy (studygurl14):

@25Mattman

OpenStudy (anonymous):

never mind, sorry @StudyGurl14

OpenStudy (studygurl14):

lol ok